OpenStudy (anonymous):

Solve: log(base c)81=4

OpenStudy (anonymous):

\[\log_c 81=4\iff c^4=81\iff c^4=3^4\iff c=3\]since we are talking about logs here, -3 is omitted as an answer, even though (-3)^4 also equals 81.